ou are testing the null hypothesis that there is no relationship between two​ variables, X and Y. From your sample of you n=20 determine that SSR= 80 and SSE=20.

What is the value of FSTAT?

At the a= 0.05level of significance, what is the critical value?

Compute the correlation coefficient by first computing r^2 and assuming the b1 is negative.

r^2=

r=

Find the test statistic

tSTAT=

Determine the critical values of t.

The lower critical value is

The upper critical value is
Answer #1

FSTAT =SSR/(SSE/(n-2))=80/(20/18)=72

At the a= 0.05level of significance and 1,18 degree of freedom , critical value =4.414

r2 =SSR/(SSR+SST) =80/(80+20 )=0.8

r =-√r2 = -0.894

test statistic =- √F = -√72 = -8.485

for 0.05level of significance and 18 degree of freedom ,

lower critical value is = -2.101

upper critical value is = 2.101